OpenStudy (anonymous):

Solve the equation. If necessary, round your answer to the nearest tenth. 9^2=49 A. 2.3, -2.3 B. 5.4, -5.4 C. 7,-7 D. 21, -21 Help please? Will give a medal =)

OpenStudy (bloomlocke367):

is there something missing in that equation?...

OpenStudy (anonymous):

i am not 100 percent sure but i think it is A

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

Did you mean \[x^2=49\]

OpenStudy (bloomlocke367):

and @Destinyrules1234 don't give direct answers...

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhh srry

OpenStudy (bloomlocke367):

@bvblover83014 what's your actual equation?

OpenStudy (bloomlocke367):

did you mean\[x^2=49\]

OpenStudy (anonymous):

\[9x ^{2}=49\]

OpenStudy (bloomlocke367):

ohhh. okay. first divide both sides by 9. \[49\div9=?\]

OpenStudy (anonymous):

5.4

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

yes! now take the square root of both sides. \[\sqrt{5.4}=\pm?\]

OpenStudy (anonymous):

2.32

OpenStudy (bloomlocke367):

yes! remember, though, it positive and negative 2.32. so, what's your answer?

OpenStudy (anonymous):

A

OpenStudy (bloomlocke367):

there you go! good job!! XD